### v1

# Tailored Emotional LLM-Supporter: Enhancing Cultural Sensitivity

Abstract
Large language models (LLMs) show promise in offering emotional support and generating empathetic responses for individuals in distress, but their ability to deliver culturally sensitive support remains underexplored due to lack of resources. In this work, we introduce CultureCare, the first dataset designed for this task, spanning four cultures and including 1729 distress messages, 1523 cultural signals, and 1041 support strategies with fine-grained emotional and cultural annotations. Leveraging CultureCare, we (i) develop and test four adaptation strategies for guiding three state-of-the-art LLMs toward culturally sensitive responses; (ii) conduct comprehensive evaluations using LLM judges, in-culture human annotators, and clinical psychologists; (iii) show that adapted LLMs outperform anonymous online peer responses, and that simple cultural role-play is insufficient for cultural sensitivity; and (iv) explore the application of LLMs in clinical training, where experts highlight their potential in fostering cultural competence in future therapists.111Github: https://github.com/UKPLab/arxiv2025-culturecare.
Content Warning: This paper includes examples that some readers may find offensive or triggering. These instances are presented for research purposes only and do not represent the views of the authors or affiliated institutions.
1 Introduction
Large language models (LLMs) have shown growing potential in offering emotional support, with recent work demonstrating that LLMs can provide empathetic, contextually relevant responses for individuals experiencing distress Zheng et al. (2024); Zhan et al. (2024); Ye et al. (2025). This emerging capability is particularly promising in online spaces, where peer support communities play a vital role in helping individuals navigate emotional challenges. However, culture shapes human emotional experiences and influences the stressors people encounter in daily life Markus and Kitayama (1991); Mesquita and Frijda (1992); Chun et al. (2006); Mesquita et al. (2017). As a result, one critical dimension remains underexplored — the cultural sensitivity of these LLMs’ support responses.
Effective emotional support is deeply shaped by cultural knowledge, unspoken assumptions, norms, and values that influence how distress is expressed and how support is received Taylor et al. (2007); Matsumoto et al. (2008); Kim et al. (2008). However, when LLMs are used as tools for emotional support, they may struggle to recognize these culturally embedded cues of distress in the first place Aleem et al. (2024). Even well-intentioned responses can cause harm or alienation without cultural grounding and sensitivity, rather than providing the support users seek Lissak et al. (2024); Moore et al. (2025). For example, in a collectivist culture, an LLM advising a user to cut ties with their family in pursuit of personal happiness, without acknowledging the cultural weight of familial duty, could come across as offensive or morally unacceptable. Similarly, in cultures where mental health is heavily stigmatized, a blunt recommendation to “seek therapy immediately” might intensify feelings of shame or social isolation rather than offering relief.
Recent research tries to address this issue through a focused case study on Pakistani culture Aleem et al. (2024). However, the study offers limited generalizability due to its narrow set of manually created distress scenarios, leaving the broader challenge of culturally aware emotional support largely unexplored. The lack of suitable datasets and evaluation of adaptation methods has hindered progress in this area. Therefore, in this work, we address the research questions by presenting a comprehensive multi-cultural investigation into LLMs’ ability to provide culturally sensitive emotional support.
To address this data gap, we introduce CultureCare. To the best of our knowledge, this is the first dataset designed to support the study of culturally-aware peer emotional support (Figure 1). CultureCare spans four distinct global cultures, namely Arabic, Chinese, German, and Jewish, and fine-grained annotated for both emotional support strategies and culturally-relevant signals, as shown in Table 1. The dataset consists of distress-response pairs sourced from real-world support interactions from Reddit, annotated for support type and cultural signals, allowing for nuanced development and evaluations for adapting LLMs’ responses across cultural contexts. Using CultureCare, we evaluate three state-of-the-art LLMs with tailored prompting strategies for adaptation. Through both automated evaluations and in-culture human evaluations, we find that while incorporating basic cultural information helps, a more effective adaptation requires detailed guidelines and attention to contextualized, explicit cultural signals. These LLM-generated responses are often better than human responses online.
While our primary focus is on peer support — where the LLM acts as a supportive peer rather than providing professional help — we also explore its potential in clinical training settings, namely in training psychology students for culturally sensitive therapy. Expert feedback highlights strong safety and promising utility.
To sum up, our contributions are: 1. We release the first dataset — CultureCare — for evaluating and training LLMs in culturally-aware emotional support, spanning four cultures with fine-grained annotations. The dataset comprises 1729 annotated distress messages, 1523 cultural signals, and 1041 support strategies. 2. We develop and test four adaptation strategies to guide three popular state-of-the-art LLMs toward generating culturally-sensitive support responses. 3. We provide comprehensive evaluations involving a LLM judge, in-culture human evaluators, and clinical psychologists to assess both emotional and cultural quality of generated responses.
2 Related Work
Culturally adapted LLMs. Recent research has found that LLMs predominantly reflect the perspectives of WEIRD (Western, Educated, Industrialized, Rich, and Democratic, Henrich et al. 2010) populations without any adaptation Atari et al. (2023); Johnson et al. (2022). Several studies attempted to address this issue, focusing on diverse tasks such as value alignment AlKhamissi et al. (2024); Liu et al. (2025b) or hate speech classification Zhou et al. (2023); Li et al. (2024); Adilazuarda et al. (2025). AlKhamissi et al. (2024); Tao et al. (2024) demonstrated that prompting LLMs with cultural and persona-specific information can effectively align models with diverse cultural values. However, existing work has not examined the effectiveness of these prompting methods for culturally aware emotional support — an important gap this study addresses.
Culture and mental health. Culturally sensitive counselling is a well-established consideration in clinical psychology and healthcare settings (Bernal et al., 1995; Resnicow et al., 1999; Kreuter and McClure, 2004; Taylor et al., 2007; Tao et al., 2015, among others). Prior research has explored various aspects of incorporating cultural sensitivity in practical domains outside of NLP, including the importance of cultural humility in improving therapy outcomes Owen et al. (2016), disparities in engagement and follow-up care across demographics Zeber et al. (2017), and the need to embed cultural competence in training programs Benuto et al. (2018). However, the application of LLMs in culturally sensitive mental health remains limited. Focusing on formal therapy settings, recent work Abbasi et al. (2025); Kim et al. (2025) explores LLM-generated synthetic multi-turn clinical conversations in multilingual contexts. Their focus on clinical therapy and synthetic data generation differs from ours, which centers on peer emotional support and adaptation strategies.
LLMs for emotional support. Existing work has shown that LLMs can provide empathetic and supportive responses when appropriately guided Zhan et al. (2024). Studies such as Liu et al. (2021); Zheng et al. (2024); Zhang et al. (2024) examine how LLMs respond to distress messages using various support strategies. However, they largely overlook the influence of cultural context in shaping emotional needs and support preferences. While Aleem et al. (2024) considers cultural context, its focus on a small set of scenarios from Pakistani culture limits its generalizability.
Furthermore, recent studies on LLMs in cognitive behavioural therapy Goel et al. (2025); Zhou et al. (2025); Zhang et al. (2025) emphasize clinical settings with structured treatment frameworks, offering little attention to cultural nuance. In contrast, our work focuses on peer-to-peer emotional support, where cultural understanding is essential for effective and supportive communication. We address this gap by incorporating cultural signals from diverse cultural communities into our annotation process, and by assessing emotional support responses from humans and LLMs, based not only on empathy but also on their cultural relevance.
3 CultureCare
We present CultureCare, a multi-cultural dataset with fine-grained span-level annotation of cultural signals and emotional distress.222We deliberately chose not to use language as the defining boundary of culture, recognizing that culturally influenced distress can be expressed in any language in online communities. As a result, CultureCare includes posts both in English and in the native languages associated with each culture. The language distribution varies by culture; see Appendix A.5. An overview of CultureCare’s construction pipeline is in Figure 2. Briefly, we begin by collecting culturally relevant mental health data from Reddit using targeted search queries. We then apply a combination of rule-based and LLM-based filters to remove noise. Finally, in-culture annotators and reviewers select the spans and label the data for emotional distress, cultural signals, and support strategies. The annotation categories for each dimension are detailed in Appendix A.4.
Table 1 compares CultureCare with existing resources. CultureCare is distinctive in its focus on multiple cultures, the explicit annotation of cultural signals and emotional distress.
| Real? | Cultures | Size | Annotations | Eval. Aspects | |
| Liu et al. (2021, ESConv) | ✓ | — | 1053 | DM, SS, E | E |
| Zheng et al. (2024, ExTES) | ✗ | — | 11177† | DM, SS | E |
| Zhang et al. (2024, FEEL) | Mix | — | 200 | DM, SS | E |
| Aleem et al. (2024) | ✗ | Pakistani | 7 | — | C |
| CultureCare | ✓ | Arabic, Chinese, German, Jewish | 462 | DM, SS, Cultural Signals | E, C |
| Category | AR | CH | GE | JE | All |
| posts | 110 | 141 | 119 | 92 | 462 |
| responses | 104 | 126 | 100 | 88 | 418 |
| distress messages | 397 | 399 | 402 | 531 | 1729 |
| cultural signals | 346 | 315 | 338 | 524 | 1523 |
| support strategies | 259 | 242 | 194 | 346 | 1041 |
| demographic info | 226 | 301 | 268 | 131 | 926 |
| Avg. post length | 316 | 492 | 690 | 389 | 480 |
| Avg. response length | 101 | 80 | 125 | 88 | 98 |
3.1 Data Collection and Filtering
Candidate posts. We use Reddit as our dataset source due to its global user base and peer-driven mental health discussions. Our data is collected via the Reddit API333https://www.reddit.com/dev/api/
Data collection and annotations are approved by the ethics board.. We focus on two setups: 1. searching mental health subreddits with culture-specific keywords, and 2. searching culture-specific subreddits with mental health keywords. The list of subreddits and keywords is in Appendix A.2. For each relevant post, we fetch the top-voted comment as the ideal supportive response, as it typically offers emotional support and resonates with readers, reflected in its high upvotes.
Filtering. Our initial dataset contained 9160 posts, many of which were noisy — e.g., general mental health tips, reactions to global events, or content unrelated to the target cultures. Since manual review of all posts was infeasible, we first applied rule-based filters to remove both explicit noise (e.g., URLs-only posts) and LLM-based filtering to remove implicit noise (e.g., posts lacking personal distress). This leaves us with 2671 posts for manual review. The full set of filters is detailed in Appendix A.3. After the final manual review, we retained 462 high-quality posts that clearly express culturally-situated emotional distress, with optional responses by humans. While small, this dataset allows for focused, fine-grained annotation.
3.2 Annotations and Quality Assurance
Each post was annotated by a in-culture annotator using detailed guidelines (Appendix A.4) identifying spans with emotional distress, cultural signals, and emotional support. On average, each post took approximately 10 minutes to annotate, as they were often lengthy and required attentive reading. To ensure high quality, a second in-culture annotator reviewed the initial annotations and provided comments in cases of disagreement; however, such instances were rare and typically involved additions rather than corrections.
3.3 Analysis of CultureCare
Statistics. CultureCare includes 462 posts (90% with responses444The rest 10% had deleted comments or no comments.), containing 1729 annotated distress messages, 1523 cultural signals, and 1041 support strategies. We also extract demographic details (e.g., age, gender, religion) from the posts using an LLM, yielding 926 signals. Example annotated posts are in Appendix F.
Prevalent norms and values. Appendix A.7 illustrates how culture shapes emotional distress, highlighting recurring themes such as family control, mental health invalidation, and gender expectations. Some experiences are culture-specific, including LGBTQ+ rejection in Arabic contexts and financial stress in German ones. While emotional struggles are universal, their expression and root causes depend on the culture.
Cultural signals types and support strategies. Arabic posts emphasize values and explicitly state demographic details, while Chinese and German prioritize norms and morals. Jewish posts focus heavily on concepts, like rabbi. Artifacts, knowledge, and language have a minimal presence in all cultures. In human responses, providing suggestions is most common across all cultures. We also observe that while Arabic and Chinese responses emphasize self-disclosure and affirmation, German responses are more balanced, and Jewish responses more often provide information. Asking questions is the least common strategy.
Demographic diversity. We extract detailed demographic information, revealing various patterns between age, gender, and other attributes.555In some cases, only high-level cultural indicators, such as “Arabic” or “Chinese”, are extractable. For example, extractable Jewish posts are predominantly female (80%), while extractable Arabic and Chinese posts have a high proportion of young adults (47% and 50%, respectively). The extractions also reflect diversity in birthplace, residence, marital status, education level, and profession. These variations highlight the rich cultural and social heterogeneity of the dataset. More details are provided in Appendix A.6.
Intensity and empathy scores. We additionally provide analysis on cultural differences in distress intensity and response empathy. In our data, in general, Jewish posts show the highest intensity, followed by German and Arabic, while Arabic responses demonstrate the highest empathy. See details in Appendix A.9.
4 Adaptation Methods
We experiment with basic adaptation strategies using prompts and a compound strategy by combining all strategies, prompts in Appendix C.
Standard (no suffix). By default, we prompt the model to be a Redditor, matching the context of the data. This variation serves as a baseline for comparing adaptation strategies.
Culture-informed role-playing (+culture). Building on prior research AlKhamissi et al. (2024); Tao et al. (2024), instructing LLM to role-play the cultural background of the person is a simple yet effective method for aligning LLM responses with culturally relevant values. Hence, this could enable more empathetic, appropriate responses, removing the cultural difference barrier in empathy Cikara et al. (2014); Davis (2018).
Guided principles / constitutions (+guided). Here, we provide guidelines based on CCCI-R (Cross-Cultural Counselling Inventory—Revised; LaFromboise et al. 1991a, b, see Appendix G for details), one of the widely established cross-cultural counselling competency measurements by APA (American Psychological Association). This approach aims to emulate some fundamental competency of professional counselling in terms of cultural sensitivity and awareness in an implicit cross-cultural setting.
Explicit cultural signals (+annotation). We explicitly add the data annotations of posts from our dataset to the prompt. The goal of this strategy is to understand whether explicitly providing LLMs with richer contextual information can improve understanding and output in an implicit cross-cultural setting.
Combined (+cga). In this method, we combined the above three basic strategies, culture, guided, and annotation (Figure 11, Appendix C). We modified the guidelines in the +guided strategy by removing CCCI-R items that focus on cross-cultural differences. Here, the LLM will be provided with explicit information and guidelines, as well as role-playing a person from the same culture.
5 Experimental Setup
In this work, we focus our experiment on the open-sourced LLMs. We evaluate models that are in the 7B-8B parameters range due to their popularity and performance, including: Llama-3.1-8B Touvron et al. (2023); Dubey et al. (2024), Qwen-2.5-7B Team (2024), and Aya-Expanse-8B Dang et al. (2024). Here, we prioritize open-source model adaptations for scientific reproducibility. We use default configurations for generation in our experiments, and all the adaptation strategies are implemented as system prompts.
| Model | Arabic | Chinese | German | Jewish | Average | ||||||
| Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | All | |
| Aya-Expanse-8B | 4.67 | 3.51 | 4.43 | 3.46 | 4.62 | 2.54 | 4.61 | 3.99 | 4.58 | 3.37 | 3.98 |
| +culture | 4.55 | 3.56 | 4.48 | 3.63 | 4.61 | 2.75 | 4.67 | 4.12 | 4.58 | 3.51 | 4.05 |
| +guided | 4.75 | 3.80 | 4.73 | 4.08 | 4.77 | 2.72 | 4.79 | 4.16 | 4.76 | 3.69 | 4.23 |
| +annotation | 4.77 | 3.73 | 4.77 | 3.78 | 4.80 | 2.75 | 4.81 | 4.10 | 4.79 | 3.59 | 4.19 |
| +cga | 4.84 | 4.39 | 4.82 | 4.25 | 4.84 | 3.44 | 4.91 | 4.55 | 4.85 | 4.16 | 4.51 |
| Qwen-2.5-7B | 4.16 | 3.02 | 4.26 | 3.20 | 3.89 | 2.66 | 4.27 | 3.60 | 4.15 | 3.12 | 3.63 |
| +culture | 4.05 | 3.23 | 4.05 | 3.38 | 3.91 | 2.73 | 4.28 | 3.71 | 4.07 | 3.26 | 3.67 |
| +guided | 4.41 | 3.29 | 4.49 | 3.42 | 4.32 | 2.78 | 4.54 | 3.73 | 4.44 | 3.30 | 3.87 |
| +annotation | 4.40 | 3.40 | 4.28 | 3.46 | 4.29 | 2.82 | 4.46 | 3.69 | 4.36 | 3.34 | 3.85 |
| +cga | 4.11 | 3.70 | 4.24 | 3.81 | 3.86 | 2.67 | 4.50 | 3.95 | 4.18 | 3.53 | 3.85 |
| Llama-3.1-8B | 3.79 | 2.98 | 4.20 | 3.41 | 3.81 | 2.67 | 4.22 | 3.89 | 4.00 | 3.24 | 3.62 |
| +culture | 3.75 | 3.65 | 4.11 | 3.99 | 3.74 | 2.62 | 4.34 | 4.15 | 3.99 | 3.61 | 3.80 |
| +guided | 4.22 | 3.40 | 4.57 | 3.89 | 4.26 | 2.54 | 4.59 | 4.11 | 4.41 | 3.49 | 3.95 |
| +annotation | 4.14 | 3.52 | 4.54 | 3.60 | 4.23 | 2.66 | 4.48 | 3.99 | 4.35 | 3.44 | 3.89 |
| +cga | 4.13 | 3.93 | 4.48 | 4.18 | 3.98 | 2.58 | 4.64 | 4.38 | 4.31 | 3.77 | 4.04 |
5.1 Automatic Evaluations
We perform fine-grained automatic evaluations based on both emotional support and cultural awareness. We use GPT-o3-mini OpenAI (2025) as the LLM judge due to its high ability for reasoning and correlation with human judges Tan et al. (2025); Gu et al. (2024).
Emotional supportiveness measures the basic requirements of an effective supporting message. Based on evaluation criteria from prior research Rashkin et al. (2019); Liu et al. (2021), we included the following criteria: 1. Empathy - the response should demonstrate a genuine understanding of the author’s emotions and convey timely, appropriate concern; and 2. Helpfulness - the response offers effective advice and tailored, actionable steps.
Cultural awareness measures the awareness and sensitivity of a response concerning cultural aspects. To match the desirable culturally sensitive support in LaFromboise et al. (1991a), three criteria are used: 1. Socio-political influence - the response demonstrates an understanding of the current sociopolitical system and its impact on the author of the post; 2. Knowledge - the response reflects knowledge about the target culture; and 3. Cultural context - the response perceives problem within the appropriate cultural context.
5.2 Human Evaluations
To further assess the emotional supportiveness and cultural awareness of the responses, we conducted two human evaluations: 1. Crowd evaluations with individuals from the corresponding culture; 2. Expert evaluations to assess the responses for safety and potential usefulness for professionals.
Crowd evaluation. To compare the effectiveness of different strategies, we conducted in-culture crowd-sourced evaluations using Prolific.666https://www.prolific.com/ We recruit two people per culture who are fluent in both English and the matching language of the culture.777Most Chinese cultural data comes from individuals in English-speaking countries who identify with Chinese culture, shaping their distress experiences. We selected crowd evaluators to match this context.
To evaluate the best adaptation strategy, we sampled 30 data points per culture for all 3 examined models (a total of 2160 response evaluations), along with the human response and the model-generated responses for all strategies. For each culture, we ask two evaluators to pick the best response in terms of both emotional supportiveness and cultural awareness, and aggregate their ranks of adaptation strategies.
To evaluate the best model, we sampled another 20 data points from one of the top strategies across all models we used, and recruited three annotators per culture for the best-performing model evaluation. Here, we use the same criteria as the previous evaluation. The labels are decided based on a majority vote. If there is a tie, another in-culture annotator is involved to make the judgment.888This only happens 5% for cultural awareness, 13.33% for emotional supportiveness, indicating that cultural awareness is a less subjective task than emotional supportiveness.
The instructions were given in both English and the native language of the culture. The detailed instructions are in the Appendix E.
Expert evaluation. We collaborated with two psychologists experienced in online emotional support to 1. validate the safety of the adaptation strategies, and 2. assess the utility of adapted LLMs for training psychology students in cross-cultural therapy Benuto et al. (2018). Together, we developed evaluation guidelines (Appendix H.1). For safety, 100 responses (50 by human, 50 by LLMs) were evaluated, with 15 double-annotated for agreement due to the time-intensive nature of assessment. For utility, 100 LLM-generated responses were evaluated. Agreement results are in Appendix H.2.
6 Results & Discussions
Table 3 presents the automatic evaluation results of emotion supportiveness and cultural awareness metrics, whereas Table 14 (Appendix D) displays the results for overall quality. In general, all models can respond with high fluency and appropriateness.
Showing in Table 3, the winning strategy is +cga by combining culture-informed role-playing, guidelines and explicit cultural signal annotations. +cga consistently outperforms other strategies across different cultures and models.
Compared to standard role-playing as a Redditor (no suffix in Table 3), simple culture-informed role-playing (+culture) does not improve emotional supportiveness consistently (7/12 times scores lower). In fact, +culture performs worse than providing explicit, detailed cultural signals or guidelines aimed at cross-cultural consultation for emotional supportiveness. Overall, the +guided strategy is more prominent in providing winning emotional support; however, the difference between +annotation and +guided is small.
By explicitly incorporating cultural considerations when offering emotional support (e.g., through +guided, +annotation or +cga), models generate better responses compared to +culture in both evaluation dimensions. We also have a similar observation over human evaluation results (details in §6.1).
Our results here highlight simple, culture-informed role-playing of a generic person from a culture is likely insufficient to provide context-dependent, culturally sensitive tasks; task-specific guidance or guiding model’s attention with explicit signals within the context could offer better help.
6.1 Human (Crowd) Evaluation
Adapted LLMs can outperform humans in providing responses. Overall, our in-culture evaluators select +cga and +annotation as the best strategies across models and cultures (Table 15). Interestingly, the human response is often ranked as the worst or second-worst strategy, 10 of 12 times for emotional supportiveness. This result further indicates that LLMs could provide better non-professional emotional support in online communities, as the average online human responses can be non-helpful, and fail to recognize and address emotional needs.
Human and LLM judge show moderate to strong correlations. We compute the Kendall rank correlation coefficient () between humans’ and LLMs’ ranking of adaptation strategies (five strategies). The detailed results are in Table 16 (Appendix E.2). In general, human-model correlations are moderate to strong for emotional awareness, except for German culture and Qwen’s responses for Jewish culture. Correlations for cultural awareness are weaker and even negative for German culture, despite it being a high-resource culture.
Notably, humans and the LLM judge diverge the most on Llama’s responses for cultural awareness in Kendall’s . However, as shown in Table 15 (Appendix E.2), humans and the LLM judge often agree on the best strategy, suggesting that divergences mainly stem from lower-ranked strategies. This highlights the need to improve LLMs as judges of cultural awareness.
No single winning model for all cultures. In-culture evaluators evaluate 20 responses from the strategies of each model. Aya-Expanse-8B leads in cultural awareness across Arabic, Chinese, and German contexts. In emotional awareness, Llama-3.1-8B performs best for Chinese culture, while Aya-Expanse-8B receives the highest ratings from evaluators from Arabic and German cultures. Jewish evaluators prefer Llama-3.1-8B for both cultural awareness and emotional supportiveness.
Our evaluation shows: 1. human-model agrees the majority of the time on the best adaptation strategy; 2. LLM judges show room for improvement in cultural awareness; 3. no single model excels at emotional support across cultures.
6.2 Additional Analysis
We qualitatively evaluate responses generated by Aya-Expanse-8B, which was preferred by most cultural groups. To analyze the types of emotional support provided, we use Llama-3.1-70B to categorize responses based on the framework from Liu et al. (2021). Figure 3 shows the distribution of emotional support message categories of Aya-Expanse-8B’s responses across cultures for the strategy. Additional histograms for other adaptation strategies are included in Appendix I. Overall, model-generated responses use a mix of reflection of feelings, affirmation, suggestions, and information, which is distinctly different from human support messages online that primarily offer suggestions. This pattern is consistent across cultures and adaptation strategies.
6.3 Evaluation with Clinical Psychologists
Our quantitative results and discussion are in Appendix H.3; key highlights include:
LLMs maintain balanced tone, outperform human responses for cognitive empathy, and emotional empathy. Overall, experts indicate that responses were neither overly positive (only 2%) nor overly negative (0%) compared to human responses (6% at both extremes). LLMs outperform humans in cognitive empathy (30% versus 4%) and emotional empathy (34% versus 2%) with the “Strongly agree” rating.
LLMs do not escalate nor introduce new emotional distress. Psychologists agree that LLMs’ responses do not introduce new distress or escalate existing negative feelings with aggregated positive rating (“Strongly agree” and “Agree”) 80% of the time, compared to 28% for human responses.
Suitable for training new clinical psychologists. A total of 88% of LLM responses were found to be promising for training psychology students in cross-cultural therapy, 46% were rated as strong examples suitable for direct use, while 42% required minor adjustments.
Open-feedback from experts. Several qualitative observations emerged from the evaluation process. First, while LLMs often generated structured responses that explicitly included culturally sensitive elements, resulting in higher scores across all criteria, human responses are more authentic and spontaneous. This raises key questions about whether users might value emotionally genuine responses over polished yet generic ones in a certain context. Second, the consistently empathetic tone of LLM (e.g., “You are not alone”) raises an ethical question: should users be informed they are interacting with AI rather than a human? Third, LLMs show a strong tendency to affirm users’ perspectives, but might occasionally miss the clinical red flag, highlighting the need for caution in relying on LLMs for emotional support and emphasizing the importance of incorporating clinical judgment and ethical safeguards into system design and deployment.
7 Conclusion
The ability of LLMs to provide culturally sensitive peer emotional support has been largely overlooked in prior research. To address this gap, we introduce CultureCare, a fine-grained, multicultural dataset encompassing four cultures. Using CultureCare, we evaluate three state-of-the-art LLMs with several carefully designed adaptation strategies. Our findings reveal that: 1. Incorporating shallow cultural information is insufficient compared to using specific, contextualized, guideline-aligned cultural adaptations; 2. LLMs have shown potential to provide emotional support that, in many cases, may surpass that of anonymous online interactions. In collaboration with professional psychologists, we explore the potential of culturally adapted LLMs for training psychology students in culturally informed therapy. Experts highlight their potential usefulness. Overall, our results demonstrate the promise of culturally adapted LLMs as effective tools for enhancing emotional support and advancing culturally sensitive counselling training.
8 Limitations
In this paper, we investigate culturally aware emotional support using data from Reddit. We acknowledge that, the platform, its users and annotations may introduce representational biases, and thus do not offer a comprehensive representation of any particular culture. In future work, we aim to collect a more diverse and representative dataset through alternative sources and extensive large scale annotations.
Our work explicitly focuses on single-turn support as we focus on online peer settings (which is mostly single-turn in nature), and as a first step in validating LLMs’ ability in providing culturally sensitive support. Future work can utilize CultureCare and extend to multi-turn in different settings.
We also attempt to move away from the common practice of using language (which is often determined by data availability) or nationality (which is typically unavailable in anonymous online communities) as a boundary for cultures. While our approach has certain limitations, it also offers a novel contribution by focusing on self-identified cultural identity and on the underlying causes of emotional stress due to cultural factors as expressed by the users themselves.
Our experiments reveal that the generated responses are sometimes easily identifiable as machine-generated due to distinctive patterns and stylistic features (we refer to them as the AI voice). Prior work has highlighted potential safety concerns associated with anthropomorphic AI systems Deshpande et al. (2023); Akbulut et al. (2024); Moore et al. (2025). This raises a safety versus utility design dilemma regarding whether LLMs intended for emotional support should strive to resemble natural human communication more closely or maintain a distinctive AI voice, and even the design choice should be the same for different cultural groups. In this study, we focus on single-turn interactions, where we think the harm is easier to avoid. We recognize that these issues may become more pronounced in multi-turn conversations due to increasing reliance on interpreting contextual cues from earlier turns, and when the model is presented with adversarial text. These challenges are well beyond the scope of this work, and we plan to address them in future research.
Ethics Statement
The data used in this study are publicly available online, and the demographic information is voluntarily provided in the original posts on Reddit. While this work explores promising strategies for automatically providing culturally aware emotional support, we do not recommend using our method directly without human verification and large-scale robustness and safety testing. This research is approved by the ethics committee of the Technical University of Darmstadt (EK 121/2024).
Acknowledgements
This work was supported by the DYNAMIC center, which is funded by the LOEWE program of the Hessian Ministry of Science and Arts (Grant Number: LOEWE/1/16/519/03/09.001(0009)/98). This work has also been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81).
We thank Yael Bar-Shacha for her help and suggestions on our expert evaluation. We thank Thy Thy Tran, Doan Nam Long, Aishik Mandal, and Anmol Goel for their feedback on a draft of this paper.
Appendix A Additional Information about CultureCare
A.1 Data License
Our experiment code and annotations are publicly available for research evaluations only. Due to Reddit’s privacy and data policy, please contact the authors or refer to the README in our project GitHub for details on how to access the dataset texts.
A.2 Lists of Subreddits and Keywords
1.
Search culture subreddits with mental health keywords:
Subreddits:
r/arabs,r/algeria,r/bahrain,r/egypt,r/iraq,r/jordan,r/kuwait,r/lebanon,r/libya,r/morocco,r/oman,r/palestine,r/qatar,r/saudiarabia,r/somalia,r/sudan,r/syria,r/tunisia,r/uae,r/yemen,r/china,r/shanghai,r/beijing,r/germany,r/depression_de,r/de,r/german,r/Israel,r/hebrew,r/HebrewIsaelis,r/Judaism,r/Jewish -
Keywords: depression, depressed, anxiety, anxious, bipolar, autistic, sad, mental health, trauma, schizophrenia, schizophrenic, anger issues
2.
Search mental health subreddits with culture keywords:
Subreddits:
r/depression,r/mmfb,r/anxiety,r/bipolar,r/bpd,r/autism,r/schizophrenia,r/mentalhealth,r/traumatoolbox,r/socialanxiety,r/anger,r/offmychest,r/bodyacceptance -
Keywords: arab, german, chinese, jewish + translated mental health keywords (e.g., traurig for sad in German)
The restrict_sr parameter was set to false in the API configuration to prevent query filtering strictly by subreddit. This allows the retrieval of relevant posts from across Reddit when no matches are found within the targeted subreddit, increasing data coverage and flexibility.
A.3 Rule-based and LLM-based Filtering
Our rule-based filters removed 46% of the noisy instances. The filters are:
F1 - incomplete posts: the post has a title but empty content.
F2 - no personal distress message: the post does not contain a personal emotional distress message, i.e., there is no use of the words “I”, “me”, etc. Such posts, for example, discuss distress of other people or a tragic event in the news.
F3 - incorrect language: we restrict our data to English and the languages of our four cultures.
F4 - more links and numbers than text: we notice that posts with too many URLs and telephone numbers tend to provide information and pointers on seeking and undestanding mental health and does not express personal distress.
Our LLM-based filters removed another 46% of the data instances. We run the filters on Qwen-2.5-72B, The filters are:
F5 - no emotional distress detected: an LLM-based classifier to rule out a post if it does not contain a personal distress message. (an advanced version of F2).
F6 - wrong or no cultural context: we instruct an LLM to rule out a post if it was not culturally relevant for our four cultures.
A.4 Annotation Guidelines.
We developed our annotation guidelines over iterations, and the final guidelines are in Figure 4.
Posts. For emotional intensity annotation, we use a 3-point scale for simplicity: 1-light, 2-moderate, 3-high. For detailed definitions and examples, see Table 4.
For cultural signals annotation, we follow a modified version of the taxonomy of cultural elements based on Liu et al. (2025a), with 7 categories: Concepts, Knowledge, Values, Norms and Morals, Language, Artifacts, and Demographics. For detailed definitions of each category and examples, see Table 5.
Responses. For emotional responses, we annotated both support message strategies and empathy scores.
Following ESConv (Liu et al., 2021), the support message strategies span 8 categories: Questions, Restatement or Paraphrasing, Reflection of Feelings, Self-disclosure, Affirmations and Reassurance, Providing Suggestions, Information, and Others. For detailed definitions of each category and examples, see Table 6.
We used a 5-point scale (Liu et al., 2021) to annotate the empathy scores of the responses. For detailed definitions and examples, see Table 7.
| Rating | Explanation | Example |
| 1 (light) | The emotion is present but subtle, with mild expression or little emphasis. | I had a bit of a rough day. |
| 2 (moderate) | The emotion is clearly expressed, showing a noticeable impact without being overwhelming. | Today was pretty hard; I’m feeling down and could use some cheering up. |
| 3 (high) | The emotion is intense and strongly emphasized, often reflecting deep or overwhelming feelings. | I’m absolutely devastated. I can’t believe this happened, and I don’t know how to cope. |
| Category | Explanation | Example |
| Concepts | Basic units of meaning underlying objects, ideas, or beliefs. | Familismo in Mexican culture. It reflects the belief that family comes first, and extended family members are often involved in decisions and provide emotional and practical support. |
| Knowledge | Information that can be acquired through education or practical experience. | Five pillars of Islam |
| Values | Beliefs, desirable end states or behaviours ranked by relative importance that can guide evaluations of things. | Respect for elders in Chinese culture |
| Norms and Morals | Set of rules or principles that govern people’s behaviour and everyday reasoning. | “You should behave like a happy and thankful child” |
| Language | Specific use of slang, speech, dialects. | “That sounds so sick!” |
| Artifacts | Materialized items as the productions of human culture, they can be forms of art, tools, machines, etc. | The Christian cross |
| Demographics | Talking about nationality or ethnicity. | “We are a Chinese couple” |
| Strategy | Explanation | Example |
| Questions | Questions related to mental health status | How often do you feel this way? |
| Restatement or Paraphrasing | Describe the situation in other words. | Sounds like your family life is really tough. |
| Reflection of Feelings | Show empathy for the other person. Understand the other’s feeling and behaviour. | I understand how anxious you are. |
| Self-disclosure | Both have made the same experience and share the same feelings. | I feel the same way sometimes. |
| Affirmation and Reassurance | Help with the other person’s uncertainty. | You have done your best and I believe you will get it. |
| Providing Suggestions | Give some possible solutions. | Find a responsible adult that you can confide in, someone you trust. |
| Information | Provide proved and useful information that can help to increase the mental well-being. | Apparently, lots of research has found that getting enough sleep before an exam can help students perform better. |
| Others | Remaining strategies. Includes wishes, hopes, etc. | I hope your luck will change. |
| Rating | Explanation | Example |
| 1 (not empathetic at all) | Shows no interest in others’ feelings and may even dismiss or ignore them. | Everyone has problems. Just get over it. |
| 2 (slightly empathetic) | Shows minimal acknowledgement of others’ feelings, but lacks genuine concern or involvement. | Oh, that’s unfortunate. |
| 3 (moderately empathetic) | Recognizes others’ emotions and shows some level of concern, but doesn’t fully engage. | That sounds tough. I’m sorry you’re dealing with that. |
| 4 (quite empathetic) | Actively listens, responds with understanding, and expresses care for the other person’s emotions. | I can see why you feel that way. That must be really hard. |
| 5 (very empathetic) | Fully connects with and validates others’ emotions, offering deep understanding and support. | I’m here for you, and I really understand what you’re going through. Let me know how I can support you. |
A.5 Language Statistics
In this work, we deliberately chose not to use language as the defining boundary of culture, recognizing that culturally influenced distress can be expressed in any language in online communities. As a result, CultureCare includes posts both in English and in the native languages associated with each culture. The language distribution varies by culture, and Table 8 shows the language statistics.
| Culture | English Only | Native Lang. / Code-mixing |
| Arabic | 81 | 29 |
| Chinese | 131 | 10 |
| German | 19 | 100 |
| Jewish | 89 | 3 |
A.6 Demographic Information
We extract demographic information from posts using the prompt in Figure 5 with GPT-4o-Mini and manually validated the results. We follow a set of demographic attributes from the WVS Haerpfer et al. (2022), which are included in Table 9. In some cases, only high-level cultural indicators, such as “Arabic” or “Chinese” are extractable.
| Attributes |
| Settlement |
| Gender |
| Age |
| Born in |
| Marital status |
| Number of people in household |
| Education |
| Profession |
| Employment |
| Class |
| Religion |
Figure 6 shows the distributions of age and gender in the CultureCare dataset, computed over posts where this information is available (50% of the dataset). The gender distribution varies notably across cultures: for instance, Jewish (JE) posts are predominantly female (80%), while German (GE) posts are mostly male (63%). Age distributions also differ; Arabic (AR) and Chinese (CH) have a higher proportion of young adults (47% and 50%, respectively), whereas GE and JE show a more balanced spread across adults and other age groups, with JE uniquely having an equal representation (26%) in teenagers, young adults, adults, and older adults. These variations highlight cultural differences in demographic representation within the dataset. Table 11 summarizes the demographic diversity in the CultureCare dataset across the cultures. Notably, the dataset captures a wide variety of cultural backgrounds and life circumstances, from varied geographic origins and residences to diverse education levels and professions, reflecting rich cultural and social heterogeneity within each group.
A.7 Cultural Themes on Emotional Distress and Support
Table 10 highlights how individuals from different cultural backgrounds experience emotional distress through the lens of cultural norms and values, the most prevalent cultural signals in our data (definitions are in Table 5). For example, both Arabic and Chinese distress messages report strong family control and dismissal of mental health issues, while Germans often express emotional exhaustion and social isolation. LGBTQ+ rejection appears uniquely in the Arabic context, while financial stress is more specific to the German entries. Gender expectations are prevalent in Arabic, Chinese, and Jewish contexts. Religious influence also varies significantly: in Arabic and Jewish cultures, religion plays a central role, either as a perceived cause of distress or as a coping mechanism. Overall, the table reveals that while emotional struggles are universal, the way they are shaped and expressed is deeply rooted in cultural values and social expectations.
| Culture | Themes |
| Arabic | Mental Health Invalidation: mental health is just seen as a phase |
| Religion Over Mental Health: …all they say is I feel this way because I don’t pray | |
| LGBTQ+ Rejection: came out … as gay … I was not met with acceptance | |
| Gender Roles: …expects a woman to depend fully on a man and her family | |
| Strict Parenting: my dad is extremely strict | |
| Chinese | Mental Health Invalidation: Depression doesn’t exist |
| Verbal Abuse: …emotionally abuse me … how much of a failure and mistake I am | |
| Gender and Identity Issues: Asian men are seen as non-masculine … weak | |
| Family Dynamics: It feels like she only lives her life for me | |
| Strict Parenting: I live with my tiger parents | |
| German | Relationships and Emotional Dynamics: my partner needs time to process emotions |
| Mental Health and Coping: I feel exhausted from pretending to be okay | |
| Family and Childhood Trauma: my parents divorced when I was young … it broke me | |
| Social Isolation: drinking is the only way I sometimes connect socially | |
| Financial Stress: I’m on social welfare and feel ashamed | |
| Jewish | Community and Social Stigma: In my community a broken engagement is … a major embarrassment |
| Personal Spiritual Practice: ask god to help those I love and the people around me | |
| Conversion and Identity: I have a long history of being told the importance of exact halachic adherence | |
| Gender Roles: …men have more responsibilities … may not be … much fun | |
| Religion and Mental Health: I’ve started going … synagogue … meet with the rabbi for a 1-on-1 session |
| Culture | Attributes |
| Arabic | gender: [male, female], age range: [14 to 40], birth place: [Iraq, arabic gulf countries, Arab country, Lebanon, Canada, Palestine, Middle east, U.S., Saudi Arabia, Syria], residence: [France, Sudan, California, UAE, Europe, Saudi Arabia, Dubai, Canada, West Bank, Iran, Aramoun (Lebanon), Bahrain, Syria, Nablus, Egypt], education: [college, high school, homeschooler, post graduate, middle school, bachelor’s degree, master’s degree], employment status: [employed, unemployed], marital status: [single, married], household size: [2-7 people], professions: [cybersecurity analyst, software engineer, pharmacist, medical worker, soldier, family business, ..], religion: [islam, christianity, atheism] |
| Chinese | gender: [male, female], age range: [15 to 58], birth place: [China, Taiwan, Malaysia, Shanghai, New Zealand, Cambodia, U.S., Europe, Australia], residence: [Fuzhou, Xingtai, Singapore, Shanghai, Malaysia, New Zealand, Hong Kong, Paris, China, Beijing, Texas, New York, Melbourne, Tokyo, North Carolina, U.K., Boston, North America], education: [high school, PhD, college graduate, master’s degree, not completed, college junior], employment status: [employed, unemployed], marital status: [single, married, engaged, in a relationship, dating], household size: [1-4 people], professions: [graphic designer, BL artist, researcher, business expansion lead, English tutor, martial arts practitioner, florist, ..], religion: [christianity, buddhism] |
| German | gender: [male, female], age range: [16 to 42], birth place: [Germany], residence: [Germany, Berlin, Ostwestfalen, Horrem, Rheinland Pfalz, Dresden, Cologne, Kleinstadt], education: [Gesamtschule, German Abitur, vocational training, German Gymnasium, high school, Lehramtsstudium, Hauptschulabschluss, Bachelor], employment status: [employed, unemployed], marital status: [single, married, divorced, engaged], household size: [1-5 people], professions: [pizza delivery, social worker, caregiver, retail, optician, service worker, real estate agent, ..], religion: [rarely mentioned] |
| Jewish | gender: [male, female], age range: [12 to 47], birth place: [U.K., Former Soviet Union], residence: [Paris, Bay Area, New York, New Hampshire], education: [student, graduate school, college], employment status: [employed full-time, employed part-time], marital status: [married, single, engaged, separated, household size: [1-4 people], professions: [intern, project manager, artist, soldier, rabbi, nurse, ..], religion: [Judaism, Agnostic, Atheism, secular Jew, Ashkenazi Jew] |
A.8 Types of Cultural Signals and Support Strategies per Culture
Figure 7 reveals the differences between the types of cultural signals among the 4 cultures. Each culture prioritizes certain signals over others, reflecting unique cultural characteristics. Arabic posts emphasize values and tend to state culture-specific demographics more explicitly (e.g., I’m an arab guy from Sudan). Chinese and German posts prioritize norms and morals, indicating a strong focus on shared principles. In contrast, Jewish posts place a significant emphasis on cultural concepts (e.g., rabbi, Shomer Shabbat), with norms and morals playing a secondary but still important role. Across all cultures, signals of types artifacts, knowledge, and language have minimal presence.
The figure also illustrates cultural differences in emotional support strategies in human responses. The most prominent strategy in all cultures is providing suggestions. Beyond the top strategy, Arabic and Chinese, both emphasize self-disclosure and affirmation. German strategies are more balanced, and Jewish culture uses the information strategy more than the 3 other cultures. The least common strategy is asking questions.
A.9 Intensity of Emotional Distress and Empathy of Responses
Table 12 shows the average scores for the intensity of distress messages in posts and the empathy of human responses across cultures and overall. Jewish posts exhibit the highest intensity (2.07), followed closely by German posts (1.98), indicating that these cultures potentially express distress more vividly compared to Arabic and Chinese. In terms of empathy, Arabic responses show the highest level (3.27), followed by Jewish (3.18), suggesting a more explicit (verbal) approach to emotional support in these cultures. In contrast, Chinese responses demonstrate the lowest empathy (2.15), while German responses fall in the middle (2.73), reflecting a more moderate engagement. Overall, the data highlights cultural variability in both the expression of distress and empathetic responsiveness. It is important to note that since annotators evaluated posts and responses from within their own cultures, these scores may reflect culturally internalized norms. In future work, we plan to explore how perceptions shift when out-of-culture annotators assess the same content.
| Culture | Intensity | Empathy |
| AR | 1.81 | 3.27 |
| CH | 1.77 | 2.15 |
| GE | 1.98 | 2.73 |
| JE | 2.07 | 3.18 |
| All | 1.89 | 2.77 |
Appendix B Infrastructure
We generate responses using the default temperature setting for all models with full precision. All our experiments were conducted using a single Nvidia A6000 GPU.
Appendix C Prompts
C.1 Adaptation Prompts
C.2 Auto-evaluations Prompts
The evaluation prompts are outlined in Figures 13 to 20. Our evaluation covers both emotional and cultural aspects, as well as the overall response quality. Metrics are in Table 13.
| Metric | Definition |
| Empathy | Measure the frequency and depth of empathy exhibited by the response. Evaluate whether the response shows a genuine understanding of the post’s emotions and whether its responses reflect timely and appropriate concern. |
| Helpfulness | Evaluate the ability of the response to provide practical solutions and assistance during the dialogue. Consider whether the model offers effective advice and actionable steps tailored to the post’s specific problems, such as emotional distress or requests for help. |
| Socio-political influence | The responder understands the current sociopolitical system and its impact on the author of the post. |
| Knowledge | The responder demonstrates knowledge about the author of the post’s culture. |
| Cultural context | The responder perceives the problem within the appropriate cultural context of the author of the post. |
| Fluency | Is the response fluent and understandable? |
| Communication | The responder’s communication is appropriate for the author of the post. |
Appendix D Additional Automatic Evaluation Results
Table 14 shows the automatic evaluation for overall quality. Models consistently score high in all strategies. This is expected as modern LLMs exhibit language influences, especially in English and high-resource languages.
| Model | AR | CH | GE | JE |
| Aya-Expanse-8B | 4.99 | 4.99 | 4.97 | 4.98 |
| +culture | 4.94 | 4.99 | 5.00 | 4.99 |
| +guided | 5.00 | 5.00 | 4.99 | 4.99 |
| +annotation | 5.00 | 5.00 | 4.95 | 4.99 |
| +cga | 5.00 | 5.00 | 4.99 | 5.00 |
| Qwen-2.5-7B | 4.76 | 4.98 | 4.55 | 4.94 |
| +culture | 4.83 | 4.97 | 4.74 | 4.98 |
| +guided | 4.99 | 4.99 | 4.97 | 4.97 |
| +annotation | 5.00 | 5.00 | 4.92 | 4.98 |
| +cga | 4.77 | 4.99 | 4.40 | 4.99 |
| Llama-3.1-8B | 4.39 | 4.95 | 4.79 | 4.89 |
| +culture | 4.30 | 4.93 | 4.63 | 4.95 |
| +guided | 4.63 | 4.96 | 4.83 | 4.97 |
| +annotation | 4.44 | 4.99 | 4.91 | 4.97 |
| +cga | 4.62 | 5.00 | 4.84 | 4.98 |
Appendix E Crowd Evaluations
E.1 Instructions
We show our crowd evaluation instructions in Figure 21. Both best adaptation strategy evaluation and best model evaluation use the same instructions, and we also translate the instructions into a language aligning with the culture.
E.2 Additional Crowd Evaluation Results
We present additional crowd evaluation results in Table 15. Our results show that explicitly providing cultural signals ( / ) and using a combined strategy () are the most preferred strategies. Further, humans and models agree majority of the time for the best adaptation strategies (shaded cells in Table 15), even though Kendall’s might be low.
| Model | Arabic | Chinese | German | Jewish | ||||
| Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | |
| Aya-Expanse-8B | +cga | +cga | +a | +a | +a | +a | +a | +a |
| Qwen-2.5-7B | +g | +cga | +a | +cga | +cga | +cga | +a | +cga |
| Llama-3.1-8B | +cga | +cga | +g | +cga | +g | +a | +cga | +cga |
| Model | AR | CH | GE | JE |
| Aya-Expanse-8B | 0.80 | 0.60 | 0.20 | 0.60 |
| Qwen-2.5-7B | 1.00 | 0.60 | 0.00 | -0.32 |
| Llama-3.1-8B | 0.60 | 0.80 | 0.40 | 0.60 |
| Aya-Expanse-8B | 0.20 | 0.60 | -0.20 | 0.00 |
| Qwen-2.5-7B | 1.00 | 0.80 | -0.40 | 0.60 |
| Llama-3.1-8B | 0.60 | 0.00 | -0.60 | 0.20 |
Appendix F Examples
We show annotated examples from CultureCare in Tables 17 and 18. These examples show culturally nuanced mental health posts, the corresponding human, and (adapted) LLM-generated responses. Each post is annotated with colour-coded signals for distress messages and their intensity, cultural signals and their categories, and support strategies.
In the first Arabic example, the person in distress reflects on how mental health struggles are dismissed in their community, noting that it is often seen as a “phase” and met with advice to “just live with it”. This shows the role cultural beliefs play in shaping stigmatizing attitudes. Both the human and LLM responses show solidarity and cultural understanding, acknowledging the taboo around mental health in Arab cultures. The LLM response further affirms the poster’s step toward seeking help. Finally, it offers a gentle suggestion to connect with others, which respects the cultural context while encouraging progress.
Another strong case appears in the Chinese example in Table 18, where the poster details high family expectations and the cultural pressure to “achieve at any cost”. They describe a deep emotional burden, exacerbated by their parents’ denial of depression as a legitimate condition. The human response connects through shared cultural experience and offers practical advice to find a trustworthy adult. The LLM echoes cultural understanding by acknowledging how depression is often dismissed in Chinese culture, and even recommends seeking a therapist familiar with those values, illustrating how culturally aligned support can validate distress while guiding users toward constructive steps. These examples underscore the value of culturally grounded empathy in both human and AI-generated responses.
| Culture | Arabic |
| Post | in Arab culture (demographics), mental health is just seen as a phase (values).. They always tell you to just ’live with it’, and if you’re in a Muslim household read the Quran (artifacts) .. and completely dismiss your feelings (moderate). Waiting in June to get myself an assessment tho (light).. |
| Human response | ..I can relate (self-disclosure)…as a society we’re still only just starting to get over the stigma around it (values).., tell you things like “everyone deals with this, just suck it up” and consider it shameful to be seeing a therapist (restatement or paraphrasing)… |
LLM
response |
.. It’s really frustrating to deal with such a stigma around mental health, especially within our cultures where these conversations are often taboo. The fact that you’re taking the initiative to seek help is a huge step forward…Many people in similar situations have found support and understanding… Consider reaching out to local support groups or online communities where you can connect with others who share similar experiences.. |
| Culture | Chinese |
| Post | 20 y/o asian male. I’ve had depression for probably 8 years (moderate). My parents constantly pressure me to fit her ideals. I have to fulfill the expectations of being above the rest.. (values).. Both my parents are strict (norms).. |
| Human response | I know how you feel (reflection of feelings) 15 y/o male in high school. My parents demand nothing less than the best..(values) Hang on and remember that you’re not alone out there (affirmation) |
LLM
response |
.. You’re carrying a lot, and it’s okay to feel burnt out and lost. In many of our cultures, mental health isn’t taken seriously, and the pressure to meet family expectations can be intense. But your worth isn’t defined by grades or appearances… |
| Culture | Jewish |
| Post | Does anyone else feel a little lonely (light) without being with anyone else Jewish? I am studying in Paris for this semester and it is great, don’t get me wrong, but I miss my people. I miss bagels and lox, and more than anything this time has made me realize that I wish I had a Jewish girlfriend.. I can confidently say I want to make my grandparents proud by being with another jew (norms). Any advice? |
| Human response | There are tons of Jews in Paris. Try to connect with the Jewish community there. (suggestion) |
LLM
response |
I totally get it, being in a foreign country can be tough, especially when you’re surrounded by people from different cultures and backgrounds. It’s great that you’re aware of your identity and what you’re looking for in a partner. Have you considered reaching out to the local Jewish community in Paris? There are probably synagogues and Jewish organizations that have social events or gatherings where you can meet like-minded people. |
| Culture | Chinese |
| Post | ..my parents don’t understand because depression doesn’t exist in the Chinese culture they grew up in (values) And the problem is that I’m really good at seeming happy. They love me so much but also want me to fulfill their dream of the ideal child - achieve at any cost, follow the blueprint of conventional success (values).. I fake being happy (light) so well that whenever I try to share how bad the pain is (moderate) , they accuse me of manipulating them.. They constantly tear me apart (moderate), first for not being good enough and now for struggling with mental illness. It’s so, so brutal (high). |
| Human response | Yes, collectivist societies tend to scoff at depression (values) If they aren’t helping, there isn’t any use confiding in them.. Find a responsible adult that you can confide in, someone you trust. (suggestion) |
LLM
response |
I’m really sorry you’re going through this. You’re not alone. Your parents may mean well, but their words can still be hurtful. As a fellow Chinese person, I know depression isn’t always understood in our culture.. You might try explaining depression in relatable terms, and consider seeing a therapist familiar with Chinese values to help bridge the gap. Most of all, remember: .. You are strong, worthy, and deserve kindness. |
| Culture | Arabic |
| Post | ..Lebanese here (demographics)(22M). Recently came out to family as gay. Unfortunately, I was not met with acceptance or support. (moderate) They did not expect this from me, and are fearful for my future. They are afraid that I will end up alone, shunned by my whole family.. My family has become even more religious after my coming out (values) and it really has put me in a spiral of depression. (high) |
| Human response | ..I am a heterosexual cisgender man who has the luck to have a personal life that is as conventional as it gets (wife, baby…). I just comment so that hopefully you feel a bit less alone. Whatever you chose to do, I hope that you’ll find serenity in your life. (affirmation) I am a practicing Muslim, but I refuse to believe that God, so merciful, would punish you for something you didn’t chose. (values) |
LLM
response |
Coming out in a conservative Arab society is incredibly difficult, but you’re not alone. .. If your family’s reaction is hurtful, it reflects their fears and beliefs, not your value. .. there are Muslim scholars who interpret sacred texts in ways that affirm LGBTQ+ lives, grounded in values like compassion and justice. Faith and identity don’t have to be in conflict. |
| Culture | German |
| Post | ..So I have a not great relationship with my parents, especially my dad (moderate). Growing up with european (german) parents, you of course feel lots of pressure to be smart and all that (norms and morals) .. |
| Human response | ..You’re not an extension of your parents (reflection on feelings) .. |
LLM
response |
Hey, I’m really sorry to hear about the challenges you’re facing with your parents and the constant need for approval. It sounds like you’re carrying a heavy burden, and it’s understandable that you want to reclaim your mental space. .. |
Appendix G Details about CCCI-R
CCCI-R LaFromboise et al. (1991a, b) is a commonly used and well-studied 20-item questionnaire developed to assess the cross-cultural counselling competency of professional counsellors and counsellors in training. In this work, we adapted a subset of items from the CCCI-R that align with the goals of this paper, as many items in the original questionnaire focus on counsellors’ professional responsibility and in-person non-verbal communication.
Appendix H Expert Evaluation
H.1 Evaluation Guidelines
The expert evaluation guidelines are in Figure 22. We developed these guidelines in close collaboration with 2 clinical psychologists to ensure their relevance to real-world therapeutic contexts. The process involved two rounds of detailed feedback and refinement. We shared an initial draft of the evaluation criteria with the clinicians, who are also responsible for training new psychologists, and who provided comprehensive input on both the conceptual framing and the specific rating questions. Based on their suggestions, we revised the structure to better capture key aspects of therapeutic communication, such as emotional empathy, emotional regulation, and cultural sensitivity. The clinicians reviewed the updated version and confirmed that no further changes were necessary. The final version reflects this collaborative, expert-informed process and is designed to support nuanced and clinically grounded evaluations.
H.2 Inter-annotator Agreement
We show the inter-annotator agreement results in Table H.2. The original (Orig.) numbers show the agreement on our original rating scales (as shown in Figure 22). The annotators have a high agreement on what is an overly positive or an overly negative response. For the rest of the criteria, we see moderate to low agreement. After closer inspection, we notice that the clinicians, in many cases, align in attitude towards a certain response, but do not align in the level of agreement. That is, they might both think that a certain response shows some emotional empathy, but one would choose to “Strongly Agree” and the other to just “Agree”. We verify this observation by reporting grouped (Grou.) agreement numbers. We find that collapsing categories to allow for partial agreement improves the agreement levels of all metrics.
| Metric | Ov. Pos. | Ov. Neg. | Cog. Emp. | Em. Emp. | Em. Reg. | Cult. Und. | Use. for train. | |
Exact match | 0.77 | 1 | 0.46 | 0.54 | 0.31 | 0.57 | 0.43 |
| Cohen’s kappa | 0.47 | 1 | 0.11 | 0.24 | 0.09 | 0.21 | 0.18 | |
Exact match | 0.93 | 1 | 0.93 | 0.86 | 0.71 | 1 | 0.73 |
| Cohen’s kappa | 0.77 | 1 | 0.85 | 0.74 | 0.44 | 1 | 0.57 |
H.3 Results
Our numerical results from the expert evaluation are in Figure 23. Overall, LLMs consistently avoid extreme (both negative and positive) emotional tones when providing responses. The empathy and emotional regulation metrics further highlight LLM strengths. For cognitive empathy, the majority (78%) of LLM responses were rated as “Strongly agree” (30%) or “Agree” (48%). Emotional empathy showed similar results, 34% ”Strongly agree” and 46% “Agree”. Emotion regulation followed the same pattern, 36% “Strongly agree” and 44% “Agree”. Human ratings in these categories were more distributed, with lower agreement levels and higher disagreement. On cultural understanding, 67% of LLM responses were rated “yes”, 31% “partially”, and only 2% “no”. When evaluating LLM usefulness for training new psychologists, 46% were seen as strong examples, 42% as promising with needing revision, and only 12% as not suitable, suggesting LLMs offer reliable, culturally aware, and emotionally attuned outputs with clear training potential.
Appendix I Distributions of Support Message Categories
We include the distributions of support message categories in Figure 24 for all strategies from the Aya-Expanse-8B model. The strategy tends to elicit more self-disclosure responses compared to the simpler strategy, even though both attempt to role-play a person from the same culture as the author of the posts. Furthermore, elicits more support responses in the “other” category (e.g., I wish you the best of luck). Overall, variations among strategies and cultures are small.